OpenStudy (anonymous):

write an equation that expresses the relationship. then solve the equation for h. F varies directly as the cube of u and inversely as h.

OpenStudy (jdoe0001):

$$ F, u \ \ cube \ \ =u^3, h \ \ inverse \ \ =\cfrac{1}{h}\\ F = u^3 \cfrac{1}{h} \implies F = \cfrac{u^3}{h} $$